# Assessing Universal Relations for Rapidly Rotating Neutron Stars: Insights from an Interpretable Deep Learning Perspective

### v1

@@ IV.1.2 Relation for the star’s reciprocal of the normalized moment of Inertia
For each particular NS model on the test subset $X_{1}$, we investigate the influence of the parameter space values on the model’s inference. From Fig. 9 (top panel), it is evident that extremely high parameter values of $\chi$, and $\log\bar{Q}$ contribute negatively to the model’s prediction. These NS configurations generally correspond to less compact stars, characterized by smaller values of parameter $\mathcal{D}$ and consequently larger moments of normalized inertia. Conversely, significantly small values of spin feature $\chi$ primarily contribute positively to the estimator’s prediction, while the lower quadrupole $\bar{Q}$ values, associated with more compact NS models, can impact the estimator’s output either positively or negatively. The combination of features with low parameter values corresponds to the most massive NS models, characterized by smaller normalized moments of inertia $\bar{I}$.
In Fig. 9 (bottom panel), the mean absolute SHAP value for each feature across the NS models in the test subset is displayed. The corresponding mean value for the rotation parameter $\chi$ is $9.28\times 10^{-4}$, whereas the quadrupole deformation parameter $\bar{Q}$ exhibits a significantly higher value of $5.93\times 10^{-3}$. The spin parameter $\chi$ enables the model to capture rotational patterns across the NSs ensemble, while the logarithmic quadrupole deformation $\log\bar{Q}$ offers deeper insights into the star’s structure by encoding information related to both rotation and the EoS. Together, these features significantly contribute to the inference performance of the DL model (12) and its ability to recognize hidden patterns in the data, enabling accurate predictions of the normalized parameter $\mathcal{D}$ across diverse Eoss and NS configurations.
@@
Table 5 summarizes the performance measures of the regression model defined in Eq. (13) as evaluated on the test set. In this formulation, the DL model accurately reproduces the data across the test set, demonstrating improved performance over the previous model with relative deviations $\leq 2.35\%$.
@@
Figure 10: Same as Fig. 7 for the new model representation provided via Eq. (13).
@@
To further evaluate the performance of the regression model in Eq. (13), Fig. 11 illustrates the predicted $\mathcal{D}(C,\log\bar{Q})$ data across the test dataset, evaluated for an array of fixed $\chi$ values. Furthermore, the observed variation in the predictions—consistent with the underlying data distribution—stems from the model’s inference within a lower-dimensional projection of the examined quasi-universal space. As expected, the most pronounced deviations from the model occur for the most compact stellar configurations. However, across all cases, each theoretical representation shows strong agreement with the underlying data, reinforcing the quasi-universal nature of the newly presented $\mathcal{D}$$C$ correlation across the employed ensemble of EoSs.
Figure 11: Predicted $\mathcal{D}(C,\log\bar{Q})$ data from the regression model (13) for a discrete set of spin values, $\chi\in[0.2,0.7]$. In each case, the model predictions align well with the corresponding test data, further supporting the model’s reliability. The vertical color bar indicates the stellar quadrupole deformation, expressed as $\log\bar{Q}$.
In addition, in terms of SHAP values, we further examine how variations within the parameter space influence the new model’s output, leading to accurate results. As shown in the top panel of Fig. 12, extremely high values of $\chi$, $C$, and $\log\bar{Q}$ tend to reduce the model’s inference, while lower values of these parameters contribute positively. This fact aligns with the earlier findings, where high values of $\chi$ and $\log\bar{Q}$ were associated with negative contributions. The key difference here lies in the behavior of the $C$ parameter: the higher values of $C$, lead to a decrease in the output value, just as we observed for $\chi$ and $\log\bar{Q}$. In contrast, lower values of stellar compactness lead to a positive contribution.
@@
In the bottom panel of Fig. 12, the mean absolute SHAP values for each feature across the NS models in the test subset are shown. The corresponding values for the parameters $\chi$, $C$, and $\log\bar{Q}$ are $7.86\times 10^{-2}$, $1.42\times 10^{-1}$, and $1.58\times 10^{-1}$, respectively. These parameters collectively play a significant role in the performance of the suggested DL model (13), aiding in the identification of underlying patterns within the data. This, in turn, supports the model’s ability to make accurate predictions of the normalized parameter $\mathcal{D}$ across the utilized ensemble of EoSs.
@@
We utilize the initial training and validation sets for training and testing, excluding parameters that lie outside the ranges typical of relatively slowly rotating stellar configurations. The regression model (14) that best reproduces the particular test data and the corresponding relative deviations histogram are illustrated in Fig. 13 top and bottom panels, respectively.
@@
The model (14) quite well reproduces the data in the test set, showcasing satisfactory precision with relative deviation better than $1.77\%$. It should be highlighted that only $0.24\%$ of the relatively slowly rotating NS configurations exhibit relative deviations exceeding $1\%$. These data are primarily associated with less compact stars, characterized by $C\lesssim 0.14$ and $\bar{Q}\gtrsim 8.12$. Concerning the variability of errors across the examined EoS categories, it is noteworthy that the hybrid EoS category exhibits the highest deviation, whereas the hadronic and hyperonic classes display lower maximum relative deviations of about $1.69\%$ and $1.37\%$, respectively (see Appendix C for a review). Among the employed EoS categories, the highest deviations are found to be linked to the DS(CMF)-2 Hybr, GPPVA(TW), and DS(CMF)-1 EoS models.
@@
Using the test set comprising relatively “slowly” rotating NS configurations, we then proceed to analyze how different values within the parameter space influence the model’s prediction. As illustrated in Fig. 14 (top panel), both the rotation parameter $\chi$ and the compactness $C$ exhibit contributions that can either increase or decrease the model’s inference. In general, NSs that are more compact or exhibit higher rotation rates tend to have a positive impact on the estimator’s prediction, while less compact stars or those with very low spin contribute negatively.
@@
Furthermore, Fig. 14 (bottom panel) presents the average absolute SHAP values for each feature across the ensemble of relatively slowly rotating stellar models. The mean values for the dimensionless spin parameter $\chi$ and the compactness $C$ are $1.95\times 10^{-4}$ and $9.52\times 10^{-4}$, respectively. While $\chi$ supports the model in identifying rotational patterns within the data, the compactness $C$ contributes more substantially to the model’s inference, as reflected by its higher average absolute SHAP value. This reflects the established correlation between the $\bar{I}-C$ and the $\mathcal{D}-C$ relations in the slow-rotation regime. Nevertheless, both parameters play integral roles in the model’s inference process, each offering complementary physical information critical to accurately estimating the normalized spin $M\times\hat{f}$.
@@ IV.2.1 Relations for the star’s reduced Quadrupole deformation
Table 6: Evaluation measures for the parametrization given by Eq. (15) regarding the training and the test set.
@@
As further validation of the effectiveness of the regression model given in Eq. (15), Figs. 16 and 17 display the corresponding predicted data over the test dataset, evaluated for a discrete array of fixed $\chi$ values. Each $\bar{Q}(C)$ and $\bar{Q}(\mathcal{D})$ data representation that arises form the DL model demonstrates excellent agreement with the underlying test data, reinforcing the quasi-universal nature of the $\bar{Q}-C$ relation as well ass the universal connection of the $\bar{Q}-\mathcal{D}$ parameters across a broad set of EoS categories for rapidly rotating NS star sequences as long as the spin parameter $\chi$ remains fixed.
Figure 16: $\bar{Q}-C$ plane: Predicted $\bar{Q}(C,\mathcal{D})$ data from the regression model (15) for a discrete set of spin values, $\chi\in[0.2,0.7]$. In each case, the model predictions closely follow the corresponding test data, further supporting the model’s reliability. The vertical color bar indicates the reciprocal of the normalized moment of inertia, $\mathcal{D}$.
Figure 17: $\bar{Q}-\mathcal{D}$ plane: Predicted $\bar{Q}(C,\mathcal{D})$ data from the regression model (15) for a discrete set of spin values, $\chi\in[0.2,0.7]$. In each case, the model predictions accurately reproduce the corresponding test data. The vertical color bar represents the stellar compactness, $C$.
@@
Using the already defined test subsample $X_{1}$, we further investigate the influence of the model’s (15) dependent parameters on inference results. As shown in Fig. 18 (top panel), each parameter can contribute either positively or negatively to the model’s prediction. For each $\bar{Q}$ estimation, higher values of parameters $\chi$ and $\mathcal{D}$ are associated with negative contributions to the estimator’s prediction, while higher values of $C$ exert a positive impact. In contrast, less compact stellar models generally have a negative impact on the estimator’s output, while models with low dimensionless angular momentum and $\mathcal{D}$ tend to exert a positive effect.
@@
In Fig. 18 (bottom panel), the mean absolute SHAP value for each parameter of interest across the NS models in the test subset is displayed. The average values of the parameters $\chi$, $C$, and $\mathcal{D}$ are $0.19$, $0.56$, and $2.25$, respectively. While the parameters $C$ and $\chi$ help the DL model to capture the physical structure and rotational dynamics of the stellar configurations, the parameter $\mathcal{D}$ emerges as the most influential feature in determining the reduced quadrupole $\bar{Q}$, as indicated by its average SHAP value contribution. This result is anticipated, given that the reciprocal of the star’s moment of inertia and the quadrupole deformation are intrinsically connected through the star’s rotational and structural characteristics. Although the mean contribution of the $\chi$ feature appears small compared to the other parameters in the corresponding figure, it plays a crucial role in capturing the universal combination among the examined quantities. As such, its inclusion is essential for preserving the universal character of $\bar{Q}$ in the predictive model. Notably, training the suggested model using only $C$ and $\mathcal{D}$ as parameters results in a noticeable increase in the relative errors of approximately $20\%$, indicating a breakdown of universality. In all cases, the parameters $(C,\chi,\mathcal{D})$ encode essential physical properties, and their combined utilization empowers the suggested DL model in Eq. (15) to deliver accurate and robust predictions of $\bar{Q}$ across a diverse ensemble of NS models in a way that is independent of the underlying EoS.
@@ IV.2.2 Relations for the star’s reduced Octupole deformation
Figure 20: Predicted $\bar{S}_{3}(\mathcal{D},\log\bar{Q})$ data from the regression model (17), evaluated over a discrete set of spin values $\chi\in[0.2,0.7]$. In each case, the model’s predictions verify the corresponding test data, reinforcing the consistency and reliability of the learned representation. The vertical color bar encodes $\log\bar{Q}$, which reflects the star’s quadrupole deformation.
@@
Consistent with previous relations, we evaluate each NS configuration in the test subset $X_{1}$ to examine how the selected input parameters of the suggested ANN influence its predictive performance. This evaluation provides insight into the model’s ability to accurately estimate $\bar{S}_{3}$ across the defined parameter space. As illustrated in the top panel of Fig. 21, each parameter can contribute either positively or negatively to the predicted value. Higher (lower) values of $\mathcal{D}$ and $\log\bar{Q}$ generally correspond to positive (negative) SHAP contributions to the model’s output, indicating a direct correlation. In contrast, feature $\chi$ does not exhibit a clear contribution pattern, as both low and high values can influence the predicted output in either a positive or negative direction. The observed SHAP-value pattern reflects known physical dependencies of $\bar{S}_{3}$ on NS properties. More specifically, higher values of $\chi$ and $\bar{Q}$ enhance rotational deformation, leading to increased octupole moments and thus positively influencing the model’s output. In addition, larger (lower) values of $\mathcal{D}$, which correlate with higher (lower) compactness and reduced deformability, are associated with smaller (larger) quadrupole moments, thereby contributing positively (negatively) to $\bar{S}_{3}$.
@@
Finally, the bottom panel of Fig. 21 displays the mean absolute SHAP values for each input feature across the test subset. The average contributions of $\chi$, $\mathcal{D}$, and $\log\bar{Q}$ to the model’s predictions are $1.60$, $5.07$, and $8.27$, respectively.

### v2

@@ IV.1.2 Relation for the star’s reciprocal of the normalized moment of Inertia
For each particular NS configuration on the test subset $X_{1}$, we investigate the influence of the parameter space values to provide an interpretation of the model’s output. From Fig. 10 (top panel), it is evident that extremely high parameter values of $\chi$, and $\log\bar{Q}$ contribute negatively to the model’s expected value $\mathbb{E}_{X}(\hat{\mathcal{D}}_{\theta^{\star}})$. These NS configurations generally correspond to less compact stars, characterized by smaller values of parameter $\mathcal{D}$ and consequently larger moments of normalized inertia. Conversely, significantly small values of spin feature $\chi$ primarily contribute positively, while the lower quadrupole $\bar{Q}$ values, associated with more compact NS models, can impact either positively or negatively. The combination of features with low parameter values corresponds to the most massive NS models, characterized by smaller normalized moments of inertia $\bar{I}$.
In Fig. 10 (bottom panel), the mean absolute SHAP value for each feature across the NS models in the test subset is displayed. The corresponding mean value for the rotation parameter $\chi$ is $9.28\times 10^{-4}$, whereas the quadrupole deformation parameter $\bar{Q}$ exhibits a significantly higher value of $5.93\times 10^{-3}$. The spin parameter $\chi$ enables the model to capture rotational patterns across the NSs ensemble, while the logarithmic quadrupole deformation $\log\bar{Q}$ offers deeper insights into the star’s structure by encoding information related to both rotation and the EoS. Taken together, these features are informative for the DL model (14), enabling it to identify hidden patterns in the data and, therefore, produce accurate predictions of the normalized parameter $\mathcal{D}$ across a diverse ensemble of EoSs and NS configurations.
@@
Table 5 summarizes the performance measures of the regression model defined in Eq. (15) as evaluated on the test set. In this formulation, the DL model accurately reproduces the data across the test set, demonstrating improved performance over the previous model with relative deviations $\leq 2.35\%$ and consistently better evaluation measures.
@@
Figure 11: Relative error distributions on the test dataset for the proposed ANN models described by Eqs. (14) and (15).
@@
To further evaluate the performance of the regression model in Eq. (15), Fig. 12 illustrates the predicted $\mathcal{D}(C,\log\bar{Q})$ data across the test dataset, evaluated for an array of fixed $\chi$ values. Furthermore, the observed variation in the predictions—consistent with the underlying data distribution—stems from the model’s inference within a lower-dimensional projection of the examined quasi-universal space. Across all cases, each theoretical representation shows strong agreement with the underlying data, reinforcing the quasi-universal nature of the newly presented $\mathcal{D}$$C$ correlation across the employed ensemble of EoSs.
Figure 12: $\mathcal{D}-C$ plane: Predicted $\mathcal{D}(C,\log\bar{Q})$ values from the regression model (15) evaluated on the test dataset for discrete spin values $\chi\in[0.2,0.7]$. Data points are color-coded by the ANN model’s absolute relative error [%], while the red star marks the configuration with the maximum relative error. The predictions align closely with the test data, exhibiting sub-percent deviations across most of the investigated parameter space. These results highlight the model’s robustness and reliability under varying spin conditions.
In addition, using SHAP values, we provide an interpretation of how the employed parameters contribute to the model’s predicted output. As shown in the top panel of Fig. 13, extremely high values of $\chi$, $C$, and $\log\bar{Q}$ tend to reduce the model’s inference, while lower values of these parameters contribute positively. This fact aligns with the earlier findings, where high values of $\chi$ and $\log\bar{Q}$ were associated with negative contributions. The key difference here lies in the behavior of the $C$ parameter: the higher values of $C$ lead to a negative contribution in the model’s expected value, just as we observed for $\chi$ and $\log\bar{Q}$. In contrast, lower values of $C$ lead to a positive contribution.
@@
In the bottom panel of Fig. 13, the mean absolute SHAP values for each feature across the NS models in the test subset are shown. The corresponding values for the parameters $\chi$, $C$, and $\log\bar{Q}$ are $7.86\times 10^{-2}$, $1.42\times 10^{-1}$, and $1.58\times 10^{-1}$, respectively. Collectively, these parameters provide an informative feature space that the DL model (15) uses to identify underlying patterns in the data. Above all, the suggested model provides accurate predictions of the normalized parameter $\mathcal{D}$ across the utilized ensemble of EoSs and NS configurations.
@@
We utilize the initial training and validation sets for training and testing, excluding parameters that lie outside the ranges typical of relatively slowly rotating stellar configurations. The regression model (16) that best reproduces the particular test data and the corresponding relative deviations histogram are illustrated in Fig. 14 top and bottom panels, respectively. It should be noted that the apparent wireframe deformation in Fig. 14 (top panel) reflects the visualization of the highly nonlinear mapping properties of the trained ANN model across the investigated $(\chi,C)$ parameter space. More specifically, even when the model is evaluated strictly within the trained input domain (i.e., without extrapolation) on a uniform mesh grid, local regions of rapid variation appear visually warped.
@@
In any case, the model (16) quite well reproduces the data in the test set, showcasing satisfactory precision with relative deviation better than $1.77\%$. It should be highlighted that only $0.24\%$ of the relatively slowly rotating NS configurations exhibit relative deviations exceeding $1\%$. These data are primarily associated with less compact stars, characterized by $C\lesssim 0.14$ and $\bar{Q}\gtrsim 8.12$. Concerning the variability of errors across the examined EoS categories, it is noteworthy that the hybrid EoS category exhibits the highest deviation, whereas the hadronic and hyperonic classes display lower maximum relative deviations of about $1.69\%$ and $1.37\%$, respectively (see Appendix C for a review). Among the employed EoS categories, the highest deviations are found to be linked to the DS(CMF)-2 Hybr, GPPVA(TW), and DS(CMF)-1 EoS models.
@@
Using the test set comprising relatively “slowly” rotating NS configurations, we then proceed to analyze how different values within the parameter space influence the model’s prediction. As illustrated in Fig. 15 (top panel), both the rotation parameter $\chi$ and the compactness $C$ exhibit influence that can be positive or negative, indicating contributions above or below the model’s expected value $\mathbb{E}_{X}(\hat{\mathcal{F}}_{\theta^{\star}})$. In general, NSs that are more compact or exhibit higher rotation rates tend to have a positive impact on the estimator’s expected value $\mathbb{E}_{X}(\hat{\mathcal{F}}_{\theta^{\star}})$, while less compact stars or those with very low spin contribute negatively.
@@
Furthermore, Fig. 15 (bottom panel) presents the average absolute SHAP values for each feature across the ensemble of relatively slowly rotating stellar models. The mean values for the dimensionless spin parameter $\chi$ and the compactness $C$ are $1.95\times 10^{-4}$ and $9.52\times 10^{-4}$, respectively. While $\chi$ supports the model in identifying rotational patterns within the data, the compactness $C$ contributes more substantially to the model’s inference, as reflected by its higher average absolute SHAP value. At fixed $\chi$, this result supports a quasi-universal correlation between the $M\times\hat{f}$ and $C$ parameters in the slow-rotation regime, consistent with the established quasi-universal $\bar{I}-C$ and $\mathcal{D}-C$ relations. In any case, both parameters play integral roles in the model’s ability to describe the data, offering complementary physical information critical to accurately estimating the normalized spin $M\times\hat{f}$.
@@ IV.2.1 Relations for the star’s reduced Quadrupole deformation
As further validation of the effectiveness of the regression model given in Eq. (17), Figs. 17 and 18 display the corresponding predicted data over the test dataset, evaluated for a discrete array of fixed $\chi$ values. Each $\bar{Q}(C)$ and $\bar{Q}(\mathcal{D})$ data representation that arises from the DL model demonstrates excellent agreement with the underlying test data, reinforcing the quasi-universal nature of the $\bar{Q}-C$ relation as well as the universal connection of the $\bar{Q}-\mathcal{D}$ parameters across a broad set of EoS categories for rapidly rotating NS star sequences as long as the spin parameter $\chi$ remains fixed.
Figure 17: $\bar{Q}-C$ plane: Predicted $\bar{Q}(C,\mathcal{D})$ data from the regression model (17) for a discrete set of spin values, $\chi\in[0.2,0.7]$. Data points are color-coded by the ANN model’s absolute relative error [%], while the red star marks the configuration with the maximum relative error. The predictions align closely with the test data, exhibiting sub-percent deviations over most of the investigated parameter space. This demonstrates the model’s robustness and reliability across different spin conditions.
Figure 18: $\bar{Q}-\mathcal{D}$ plane: Predicted $\bar{Q}(C,\mathcal{D})$ data from the regression model (17) for a discrete set of spin values, $\chi\in[0.2,0.7]$. Data points are color-coded by the ANN model’s absolute relative error [%], while the red star marks the configuration with the maximum relative error. The predictions align closely with the test data, exhibiting deviations below one percent for most cases. This consistent performance demonstrates the model’s robustness under varying spin conditions.
@@
Using the already defined test subsample $X_{1}$, we further investigate the influence of the model’s (17) dependent parameters on the final output. As shown in Fig. 19 (top panel), each parameter can contribute either positively or negatively relative to the model’s expected value $\mathbb{E}_{X}(\hat{\bar{Q}}_{\theta^{\star}})$. For each $\bar{Q}$ estimation, higher values of parameters $\chi$ and $\mathcal{D}$ are associated with negative contributions, while higher values of $C$ exert a positive impact. In contrast, less compact stellar models generally have a negative impact, while models with low dimensionless angular momentum and $\mathcal{D}$ tend to exert a positive effect.
@@
In Fig. 19 (bottom panel), the mean absolute SHAP value for each parameter of interest across the NS models in the test subset is displayed. The average values of the parameters $\chi$, $C$, and $\mathcal{D}$ are $0.19$, $0.56$, and $2.25$, respectively. While the parameters $C$ and $\chi$ help the DL model to capture the physical structure and rotational dynamics of the stellar configurations, the parameter $\mathcal{D}$ emerges as the most influential feature, as indicated by its average SHAP value contribution. This result is anticipated, given that the reciprocal of the star’s moment of inertia and the quadrupole deformation are intrinsically connected through the star’s rotational and structural characteristics. In particular, this intrinsic connection is evident in the slow-rotation regime where the well-established $\bar{I}–\bar{Q}$ relation reflects its common dependence. Likewise, for sequences at fixed spin $\chi$, as shown previously, EoS-insensitive $\mathcal{D}-\bar{Q}$ and $\bar{I}–\bar{Q}$ relations persist (see e.g. Fig. 18 for a review). Although the mean contribution of the $\chi$ feature appears small compared to the other parameters in the corresponding figure, it plays a crucial role in capturing the universal combination among the examined quantities. As such, its inclusion is essential for preserving the universal character of $\bar{Q}$ in the predictive model. Notably, training the suggested model using only $C$ and $\mathcal{D}$ as parameters results in a noticeable increase in the relative errors of approximately $20\%$, indicating a breakdown of universality. In all cases, the parameters $(C,\chi,\mathcal{D})$ encode essential physical properties, and their combined utilization empowers the suggested DL model in Eq. (17) to deliver accurate and robust predictions of $\bar{Q}$ across a diverse ensemble of NS models in a way that is independent of the underlying EoS.
@@ IV.2.2 Relations for the star’s reduced Octupole deformation
Figure 21: $\bar{S}_{3}-\mathcal{D}$ plane: Predicted $\bar{S}_{3}(\mathcal{D},\log\bar{Q})$ data from the regression model (19), evaluated over a discrete set of spin values $\chi\in[0.2,0.7]$. Data points are color-coded by the ANN model’s absolute relative error [%], while the red star marks the configuration with the maximum relative error. In any case, the model predictions align closely with the test data, showing deviations below one percent in most of the investigated parameter space. This further indicates the model’s robustness and reliability across different conditions.
@@
Consistent with previous relations, we evaluate each NS configuration in the test subset $X_{1}$ to examine how the selected inputs are attributed to the model’s predicted output. This evaluation provides insight into the model’s ability to estimate $\bar{S}_{3}$ across the defined parameter space. As illustrated in the top panel of Fig. 22, each parameter can contribute either positively or negatively to the model’s expected value $\mathbb{E}_{X}(\hat{\bar{S}}_{3,\theta^{\star}})$. Higher (lower) values of $\mathcal{D}$ and $\log\bar{Q}$ generally correspond to positive (negative) SHAP contributions, indicating a direct correlation. In contrast, feature $\chi$ does not exhibit a clear contribution pattern, as both low and high values can influence either in a positive or negative direction. The observed SHAP-value pattern reflects known physical dependencies of $\bar{S}_{3}$ on NS properties. More specifically, irrespective of $\chi$, higher values of $\bar{Q}$ lead to increased octupole moments and thus positively influence the model’s expected value $\mathbb{E}_{X}(\hat{\bar{S}}_{3,\theta^{\star}})$. In addition, larger (lower) values of $\mathcal{D}$, which correlate with higher (lower) compactness and reduced deformability, are associated with smaller (larger) quadrupole moments, thereby contributing positively (negatively) to $\bar{S}_{3}$ expected value.
@@
Finally, the bottom panel of Fig. 22 displays the mean absolute SHAP values for each input feature across the test subset. The average contributions of $\chi$, $\mathcal{D}$, and $\log\bar{Q}$ are $1.60$, $5.07$, and $8.27$, respectively.